The manager of a crew that installs carpeting has tracked the crew’s output over the past several weeks, obtaining these figures:

Week Crew Size Yards Installed
1 3 103
2 2 72
3 4 102
4 5 58
5 6 72
6 7 53

    
a. Compute the labor productivity for each of the weeks. (Round your answers to 2 decimal places.)

Week Crew size Labor productivity
(Yards/Person)
1 3
2 2
3 4
4 5
5 6
6 7


b. Which crew size works best?

A crew of      seems to work best.
Answer #1

A. Labor productivity = Number of yards installed / crew size

Week 1 = 103 / 3 = 34.33

Week 2 = 72 / 2 = 36

Week 3 = 102 / 4 = 25.5

Week 4 = 58 / 5 = 11.6

Week 5 = 72 / 6 = 12

Week 6 = 53 / 7 = 7.57

B. A crew of 2 seems to work best.

Do give a positive rating if you think this answer helped